Solving linear equations with fractional coefficients
3/4x +1/2y=-14
-1/3x+1/6y=7

Answer by Alan3354(69443)

3/4x +1/2y=-14
-1/3x+1/6y=7
-----------
Multiply to eliminate fractions.
3/4x +1/2y=-14 times 4 --> 3x + 2y = -56
-1/3x+1/6y=7 times 6 --> -2x + y = 42
-------------
3x + 2y = -56
-2x + y = 42 times 2 --> -4x + 2y = 84
------
3x + 2y = -56
-4x + 2y = 84
------------------------- subtract
7x = -140
x = -20
==================
-2x + y = 42
40 + y = 42
y = 2